24. What is the electric flux through the surface below? Q1--1pC, Q2 3pC , Q3--ZuC. 03 CQ Q1 -Q2
Answer #1

here,

Q1 = - 1 uC = - 1 * 10^-6 C

Q2 = 3 uC = 3 * 10^-6 C

Q3 = - 2 uC = - 2 * 10^-6 C

Q4 = 1 uC = 1 * 10^-6 C

using Gauss's law

the electric flux through the surfacr below , phi = charge enclosed /e0

phi = (Q1 + Q2)/e0

phi = (-1 + 3) * 10^-6 /(8.85 * 10^-12)

phi = 2.26 * 10^5 N.m^2 /C